### v1

# Progressive Contrastive Learning with Multi-Prototype for Unsupervised Visible-Infrared Person Re-identification

Abstract
Unsupervised visible-infrared person re-identification (USVI-ReID) aims to match specified people in infrared images to visible images without annotation, and vice versa. USVI-ReID is a challenging yet under-explored task. Most existing methods address the USVI-ReID problem using cluster-based contrastive learning, which simply employs the cluster center as a representation of a person. However, the cluster center primarily focuses on shared information, overlooking disparity. To address the problem, we propose a Progressive Contrastive Learning with Multi-Prototype (PCLMP) method for USVI-ReID. In brief, we first generate the hard prototype by selecting the sample with the maximum distance from the cluster center. This hard prototype is used in the contrastive loss to emphasize disparity. Additionally, instead of rigidly aligning query images to a specific prototype, we generate the dynamic prototype by randomly picking samples within a cluster. This dynamic prototype is used to retain the natural variety of features while reducing instability in the simultaneous learning of both common and disparate information. Finally, we introduce a progressive learning strategy to gradually shift the model’s attention towards hard samples, avoiding cluster deterioration. Extensive experiments conducted on the publicly available SYSU-MM01 and RegDB datasets validate the effectiveness of the proposed method. PCLMP outperforms the existing state-of-the-art method with an average mAP improvement of 3.9%. The source codes will be released.
1 Introduction
Figure 1: Illustration for the core idea of our proposal. Left: Initial sample distribution and movement trends during optimization. Right: Sample distribution upon completion of optimization. CPCL denotes Centroid Prototype Contrastive Learning, HPCL for Hard Prototype Contrastive Learning, and DPCL for Dynamic Prototype Contrastive Learning.
Visible-infrared person re-identification (VI-ReID) aims at matching the same person captured in one modality with their counterparts in another modality Ye et al. (2021); Gong et al. (2024); Yang et al. (2024). With the development of deep learning, VI-ReID has achieved remarkable advancements Yu et al. (2023); Kim et al. (2023); Feng et al. (2023); Fang et al. (2023). However, the development of existing VI-ReID methods is still limited due to the requirement for expensive-annotated training data Pang et al. (2022). To mitigate the problem of annotating large-scale cross-modality data, some semi-supervised VI-ReID methods Wang et al. (2022); Shi et al. (2023); Yang et al. (2023b) are proposed to learn modality-invariant and identity-related discriminative representations by utilizing both labeled and unlabeled data. For this purpose, OTLA Wang et al. (2022) proposed an optimal-transport label assignment mechanism to assign pseudo-labels for unlabeled infrared images while ignoring how to calibrate noise pseudo-labels. DPIS Shi et al. (2023) integrates two pseudo-labels generated by distinct models into a hybrid pseudo-label for unlabeled infrared data, but it makes the training process more complex. Although these methods have gained promising performances, they still rely on a certain number of manual-labeled data.
USVI-ReID is gaining prominence and several methods Yang et al. (2022a); Zhang et al. (2023b); Yang et al. (2023c); Wu and Ye (2023) have proposed to tackle the issues of expensive visible-infrared annotation through contrastive learning. These methods create two modality-specific memories, one for visible features and the other for infrared features. During training, these methods consider the memory center as a centroid prototype and minimize the contrastive loss across the features of query images and centroid prototype. Then, these methods aggregate the corresponding centroid prototypes based on similarity. However, the centroid prototype only stores the common information of each person, neglecting the diverse information. Just like a normal distribution, to better reflect the data distribution of a dataset, we need not only the mean but also the variance. Additionally, contrastive loss based on the centroid prototype may blur the boundaries across different clusters as illustrated in the top row of Fig. 1.
In this paper, we rethink the design of the prototype used in contrastive learning for USVI-ReID. Firstly, we design a Hard Prototype Contrastive Learning (HPCL) to mine diverse yet meaningful information and improve the boundaries across different clusters. In contrast to traditional contrastive learning methods, we choose the hard samples to serve as the hard prototype as shown in the middle row of Fig. 1. In other words, the hard prototype is the one that is farthest from the memory center. The hard prototype encompasses distinctive information. Furthermore, we introduce the concept of Dynamic Prototype Contrastive Learning (DPCL), we randomly select samples from each cluster to serve as the dynamic prototype. DPCL effectively accounts for the intrinsic diversity within clusters, enhancing the model’s adaptability to varying data distributions. The dynamic prototype offers more comprehensive representations and relationships compared with memory samples overlapping, as shown in the bottom low of Fig. 1. Early clustering results are unreliable, and utilizing hard samples at this stage may lead to cluster degradation. Therefore, we introduce progressive contrastive learning to gradually focus on hard samples.
The main contributions are summarized as follows:
  • We propose a progressive contrastive learning with multi-prototype method for the USVI-ReID. We reconsider the design of prototypes in progressive contrastive learning to ensure that the model stably captures both commonality and diversitiy.
  • We propose Hard Prototype Contrastive Learning for mining distinctive yet significant information, and Dynamic Prototype Contrastive Learning for preserving the intrinsic diversity in sample features.
  • The experimental results on the SYSU-MM01 and RegDB datasets with superior performance against the state-of-the-art methods demonstrate the effectiveness of the proposed method.
2 Related Work
2.1 Supervised Visible-Infrared Person ReID
Visible-Infrared Person ReID (VI-ReID) has drawn much attention in recent years Ye et al. (2021); Yang et al. (2022b). Many VI-ReID works focused on mitigating huge semantic gaps across modalities have made advanced progress, which can be classified into two primary classes based on their different aligning ways: image-level alignment and feature-level alignment. The image-level alignment methods focus on reducing cross-modality gaps by modality translation. Some GAN-based methods Wang et al. (2020, 2019) are proposed to perform style transformation for aligning cross-modality images. However, the generated images unavoidably contain noise. Therefore, X-modality Li et al. (2020) and its promotion Zhang et al. (2021); Wei et al. (2021) align cross-modality images by introducing a middle modality. Mainstream feature-level alignment methods Liu et al. (2022); Zhang et al. (2022); Ye et al. (2020); Sun et al. (2022); Wu et al. (2021) focus on minimizing cross-modality gaps by finding a modality-shared feature space. To this end, SAAI Fang et al. (2023) proposes a Semantic Alignment and Affinity Inference framework, which aligns the latent semantic part features by shared prototypes, enabling the aggregation of latent semantic part features with the same semantic information across both modalities. However, the advanced performances of the above methods build on large-scale human-labeled cross-modality data, which are quite time-consuming and expensive, thus hindering the fast application of these methods in real-scenes.
Figure 2: Framework of our PCLMP. The framework consists of two stages: the first stage employs contrastive learning with centroid prototypes to learn well-discriminative representation, and the second stage introduces contrastive learning with hard and dynamic prototypes to further focus on hard samples.
2.2 Unsupervised Single-Modality Person ReID
The existing unsupervised single-modality person ReID methods can be roughly divided into two classes: Unsupervised domain adaption methods (UDA), which try to leverage the knowledge transferred from labeled source domain to improve performance Ge et al. (2020a, b), and Purely unsupervised methods (USL), which directly train a USL-ReID model on the unlabeled target domain Zhang et al. (2023b); Cho et al. (2022). Compared with the UDA methods, the USL methods are more challenging. Recently, cluster-contrast learning Dai et al. (2022) has achieved impressive performance by performing contrastive learning at the cluster level. However, cluster-contrast with a uni-proxy can be biased and confusing, which fails to accurately describe the information of a cluster. To this end, methods Zou et al. (2023); Wu et al. (2022) proposed maintaining multi-proxies for a cluster to adaptively capture different information within the cluster. The above methods are mainly proposed to solve the single-modality ReID task, but they are limited to solving the USL-VI-ReID task due to large cross-modality gaps.
2.3 Unsupervised Visible-Infrared Person ReID
Recently, unsupervised visible-infrared person ReID has attracted much attention due to the advantage of not relying on any data annotation. Some UDA methods Liang et al. (2021); Wang et al. (2022) use a well-annotated labeled source domain for pre-training to solve the USVI-ReID task. Some purely unsupervised methods Yang et al. (2022a); Wu and Ye (2023); Yang et al. (2023c, a) adopt contrastive learning to boost performance inspired by Dai et al. (2022), which mainly follow a two-step loop paradigm: generating pseudo-labels using the DBSCAN algorithm Ester et al. (1996) to create memory banks with clustering centers and establishing cross-modality correspondences based on these memory banks. However, pseudo-labels are often inaccurate and rigid, CCLNet Chen et al. (2023) leverages the text information from CLIP to afford greater semantic monitoring insights to compensate for the rigidity of pseudo-labels. Moreover, reliable cross-modality correspondences are vital to USVI-ReID, thus PGM Wu and Ye (2023) proposes a progressive graph matching framework to establish more reliable cross-modality correspondences. However, cluster centers mainly present common information while lacking distinctive information, which results in ambiguous cross-modality correspondences when meeting hard samplesYang et al. (2021, 2023d)
3 Method
3.1 Overview and Problem Formulation
We propose a Progressive Contrastive Learning with Multi-Prototype (PCLMP) method for USVI-ReID, which mainly contains an online backbone, momentum backbone, and progressive contrastive learning strategy with a centroid prototype, a hard prototype, and a dynamic prototype, as shown in Fig. 2. The online backbone is a standard network, updated through back-propagation. The momentum backbone mirrors the structure of the online backbone, updated through the weights of the online backbone. The clustering is used to generate pseudo labels for creating cluster-aware memory, and we employ DBSCAN Ester et al. (1996) for this clustering. We use progressive contrastive learning with multi-prototype to optimize the online backbone. PCLMP primarily focuses on representation learning, and we use PGM Wu and Ye (2023) to aggregate cross-modality memory.
Given a USVI-ReID dataset $D=\{V,R\}$, where $V=\{V_{i}\}^{N_{v}}_{i=1}$ represents the visible images and $R=\{R_{j}\}^{N_{r}}_{j=1}$ denotes the infrared images. $V_{i}$ and $R_{j}$ represent the set of images corresponding to the $i$-th and $j$-th class. $N_{v}$ and $N_{r}$ denote the number of visible and infrared clusters, respectively. $q^{v}$ and $q^{r}$ are visible and infrared query instance features. In the USVI-ReID task, the purpose is to train a deep neural network to obtain modality-invariant and identity-related features from images without any annotation.
3.2 Centroid Prototype Contrastive Learning
Following the USVI-ReID methods Yang et al. (2022a, 2023c); Shi et al. (2024), our method uses centroid prototype contrastive learning to optimize the online backbone $\phi_{0}$ in the first state, which includes memory initialization and optimization.
[…]
$m^{v}_{i}=\frac{1}{|V_{i}|}\sum_{v\in V_{i}}\phi_{0}(v),$ (1)
$m^{r}_{j}=\frac{1}{|R_{j}|}\sum_{r\in R_{j}}\phi_{0}(r),$ (2)
where ||$\cdot$|| ⋅ | denotes the number of instances.
Optimization. During training, we update the two modality-specific memories by a momentum updating strategy Dai et al. (2022). We treat the memory center as a centroid prototype and optimize the feature extractor $\phi_{0}$ using contrastive learning with the centroid prototype, computed as:
$L^{v}_{CPCL}=-\log\frac{\exp\left(q^{v}\cdot{m_{+}^{v}}/\tau\right)}{\sum_{i=1}^{N_{v}}\exp\left(q^{v}\cdot{m_{i}^{v}}/\tau\right)},$ (3)
$L^{r}_{CPCL}=-\log\frac{\exp\left(q^{r}\cdot{m_{+}^{r}}/\tau\right)}{\sum_{j=1}^{N_{r}}\exp\left(q^{r}\cdot{m_{j}^{r}}/\tau\right)},$ (4)
$L_{CPCL}=L^{v}_{CPCL}+L^{r}_{CPCL},$ (5)
where $m_{+}^{v(r)}$ is the positive centroid prototype, denoting a query and the prototype shares the same pseudo label. The $\tau$ is a temperature hyper-parameter.
3.3 Hard Prototype Contrastive Learning
To ensure that the prototype effectively captures both the commonality and variation in a person, we devise a novel prototype for contrastive learning, which is referred to as Hard Prototype Contrastive Learning. HPCL is designed to provide a comprehensive understanding of personal characteristics, which benefits its handling of hard samples.
We use the online backbone $\phi_{0}$ to extract feature representations, and select $k$ samples that are farthest from the memory center as the hard prototype:
$hm^{v}_{i}=\underset{{{\forall v\in V_{i}}}}{\arg\max}\left\|\phi_{0}(v)-m^{v}_{i}\right\|,$ (6)
$hm^{r}_{j}=\underset{{{\forall r\in R_{j}}}}{\arg\max}\left\|\phi_{0}(r)-m^{r}_{j}\right\|.$ (7)
Then, we construct contrastive loss with the hard prototype to minimize the distance between the query and the positive hard prototype while maximizing their discrepancy to all other cluster hard prototypes, as follows:
$L_{HPCL}^{v}=-\log\frac{\exp\left(q^{v}\cdot{hm_{+}^{v}}/\tau\right)}{\sum_{i=1}^{N_{v}}\exp\left(q^{v}\cdot{hm_{i}^{v}}/\tau\right)},$ (8)
$L_{HPCL}^{r}=-\log\frac{\exp\left(q^{r}\cdot{hm_{+}^{r}}/\tau\right)}{\sum_{j=1}^{N_{r}}\exp\left(q^{r}\cdot{hm_{j}^{r}}/\tau\right)},$ (9)
$L_{HPCL}=L_{HPCL}^{v}+L_{HPCL}^{r},$ (10)
where $hm_{+}^{v(r)}$ is the positive hard prototype representation and the $\tau$ is a temperature hyper-parameter.
Finally, we update the two modality-specific memories with a momentum-updating strategy:
$hm^{v}_{+(t)}=\alpha{hm}^{v}_{+(t-1)}+(1-\alpha)q^{v}_{t},$ (11)
$hm^{r}_{+(t)}=\alpha{hm}^{r}_{+(t-1)}+(1-\alpha)q^{r}_{t},$ (12)
where $\alpha$ is a momentum coefficient that controls the update speed of the memories. $t$ and $t-1$ refer to the current and last iteration, respectively.
The hard prototype contrastive learning has two main advantages: For intra-class feature learning, it ensures that the learning process does not just focus on the shared characteristics within a cluster but also considers the diverse elements, which are often more informative. For inter-class feature learning, it is also beneficial for increasing the distances between different persons. In contrast, centroid prototypes tend to average features, lacking diversity, which can affect the network’s ability to extract discriminative features.
3.4 Dynamic Prototype Contrastive Learning
Inspired by MoCo He et al. (2020), we design dynamic prototype contrastive learning in order to preserve the intrinsic diversity in sample features while reducing instability in the simultaneous learning of both common and diverse information. DPCL comprises an online backbone $\phi_{0}$ and a momentum backbone $\phi_{m}$. The momentum backbone mirrors the structure of the online backbone, which is updated by the accumulated weights of the online backbone:
${\phi}_{m}^{t}=\beta{\phi}_{m}^{t-1}+(1-\beta){\phi}_{0}^{t},$ (13)
where $\beta$ is a momentum coefficient that controls the update speed of the momentum backbone, which is set to 0.999. $t$ and $t-1$ refer to the current and last iteration, respectively. The momentum backbone ${\phi}_{m}$ update by the moving averaged weights, which are more stable and resistant to sudden fluctuations or noisy updates He et al. (2020).
We randomly select $M$ visible/infrared samples from each cluster, denoted as $X^{v}_{i}$ and $X^{r}_{j}$. We use the momentum backbone $\phi_{m}$ to extract feature representation and store them in visible memory $DM_{RGB}$={$F^{v}_{1}$, $F^{v}_{2}$, $\cdots$, $F^{v}_{N_{v}}$} and infrared memory $DM_{IR}$={$F^{r}_{1}$, $F^{r}_{2}$, $\cdots$, $F^{r}_{N_{r}}$}, as follows:
$F^{v}_{i}=\phi_{m}(X^{v}_{i}),$ (14)
$F^{r}_{j}=\phi_{m}(X^{r}_{j}).$ (15)
We select visible dynamic prototype $dm^{v}_{i}$ from $DM_{RGB}$. In the same cluster, we select the sample farthest from the query image as the prototype. In different clusters, we choose the sample closest to the query image as the prototype:
$dm^{v}_{i}=\begin{cases}\underset{{{\forall f^{v}_{i}\in F^{v}_{i}}}}{\arg\max}\left\|q^{v}-f^{v}_{i}\right\|&\text{if }y_{q}=y_{i}\\ \underset{{{\forall f^{v}_{i}\in F^{v}_{i}}}}{\arg\min}\left\|q^{v}-f^{v}_{i}\right\|&\text{if }y_{q}\neq y_{i}\end{cases},$ (16)
[…]
The overall optimization goal of DPCL is as follows:
$L_{DPCL}^{v}=-\log\frac{\exp\left(q^{v}\cdot{dm_{+}^{v}}/\tau\right)}{\sum_{i=1}^{N_{v}}\exp\left(q^{v}\cdot{dm_{i}^{v}}/\tau\right)},$ (17)
$L_{DPCL}^{r}=-\log\frac{\exp\left(q^{r}\cdot{dm_{+}^{r}}/\tau\right)}{\sum_{j=1}^{N_{r}}\exp\left(q^{r}\cdot{dm_{j}^{r}}/\tau\right)},$ (18)
$L_{DPCL}=L_{DPCL}^{v}+L_{DPCL}^{r},$ (19)
where $dm_{+}^{v(r)}$ is the positive dynamic prototype representation, i.e., the query image and dynamic prototype have the same identity.
DPCL promotes a flexible and adaptable learning process, aiming to minimize discrepancies between samples and their respective dynamic prototypes, rather than rigidly aligning query images with a fixed prototype.
3.5 Progressive Contrastive Learning
In the initial training phases, representations are generally of lower quality. Introducing hard samples at this period could be counterproductive, potentially leading the model optimization in an incorrect direction right from the start Zou et al. (2023); Gong et al. (2022). To address this issue, we introduce the Progressive Contrastive Learning, which forms the overall loss function:
${L}_{PCLMP}=\left\{\begin{array}[]{l}{L}_{CPCL},\quad\text{ if }\text{ epoch }\leqslant E_{\text{CPCL }}\\ \lambda{L}_{HPCL}+(1-\lambda){L}_{DPCL},\quad\text{else}\end{array}\right.$ (20)
where $\lambda$ is the loss weight, $E_{\text{CPCL}}$ is a hyper-parameter.
Table 1: Comparisons with state-of-the-art methods on SYSU-MM01 and RegDB, including SVI-ReID, SSVI-ReID and USVI-ReID methods. All methods are measured by Rank-1 (%) and mAP (%). GUR* denotes the results without camera information.
Settings SYSU-MM01 RegDB
All Search Indoor Search Visible2Thermal Thermal2Visible
Type Method Venue Rank-1 mAP Rank-1 mAP Rank-1 mAP Rank-1 mAP
SVI-ReID DDAG Ye et al. (2020) ECCV’20 54.8 53.0 61.0 68.0 69.4 63.5 68.1 61.8
AGW Ye et al. (2022) TPAMI’21 47.5 47.7 54.2 63.0 70.1 66.4 70.5 65.9
NFS Chen et al. (2021) CVPR’21 56.9 55.5 62.8 69.8 80.5 72.1 78.0 69.8
LbA Park et al. (2021) ICCV’21 55.4 54.1 58.5 66.3 74.2 67.6 72.4 65.5
CAJ Ye et al. (2021) ICCV’21 69.9 66.9 76.3 80.4 85.0 79.1 84.8 77.8
MPANet Wu et al. (2021) CVPR’21 70.6 68.2 76.7 81.0 83.7 80.9 82.8 80.7
DART Yang et al. (2022b) CVPR’22 68.7 66.3 72.5 78.2 83.6 75.7 82.0 73.8
FMCNet Zhang et al. (2022) CVPR’22 66.3 62.5 68.2 74.1 89.1 84.4 88.4 83.9
MAUM Liu et al. (2022) CVPR’22 71.7 68.8 77.0 81.9 87.9 85.1 87.0 84.3
MID Huang et al. (2022) AAAI’22 60.3 59.4 64.9 70.1 87.5 84.9 84.3 81.4
LUPI Alehdaghi et al. (2022) ECCV’22 71.1 67.6 82.4 82.7 88.0 82.7 86.8 81.3
PMT Lu et al. (2023) AAAI’23 67.5 65.0 71.7 76.5 84.8 76.7 84.2 75.1
DEEN Zhang and Wang (2023) CVPR’23 74.7 71.8 80.3 83.3 91.1 85.1 89.5 83.4
SGIEL Feng et al. (2023) CVPR’23 77.1 72.3 82.1 83.0 92.2 86.6 91.1 85.2
PartMix Kim et al. (2023) CVPR’23 77.8 74.6 81.5 84.4 85.7 82.3 84.9 82.5
ProtoHPE Zhang et al. (2023a) MM’23 71.9 70.6 77.8 81.3 88.7 83.7 88.7 82.0
CAL Wu et al. (2023) ICCV’23 74.7 71.7 79.7 83.7 94.5 88.7 93.6 87.6
MUN Yu et al. (2023) ICCV’23 76.2 73.8 79.4 82.1 95.2 87.2 91.9 85.0
SAAI Fang et al. (2023) ICCV’23 75.9 77.0 83.2 88.0 91.1 91.5 92.1 92.0
FDNM Zhang et al. (2024) arXiv’24 77.8 75.1 87.3 89.1 95.5 90.0 94.0 88.7
SSVI-ReID MAUM-50 Liu et al. (2022) CVPR’22 28.8 36.1 - - - - - -
MAUM-100 Liu et al. (2022) CVPR’22 38.5 39.2 - - - - - -
OTLA Wang et al. (2022) ECCV’22 48.2 43.9 47.4 56.8 49.9 41.8 49.6 42.8
TAA Yang et al. (2023b) TIP’23 48.8 42.3 50.1 56.0 62.2 56.0 63.8 56.5
DPIS Shi et al. (2023) ICCV’23 58.4 55.6 63.0 70.0 62.3 53.2 61.5 52.7
USVI-ReID H2H Liang et al. (2021) TIP’21 30.2 29.4 - - 23.8 18.9 - -
OTLA Wang et al. (2022) ECCV’22 29.9 27.1 29.8 38.8 32.9 29.7 32.1 28.6
ADCA Yang et al. (2022a) MM’22 45.5 42.7 50.6 59.1 67.2 64.1 68.5 63.8
NGLR Cheng et al. (2023) MM’23 50.4 47.4 53.5 61.7 85.6 76.7 82.9 75.0
MBCCM He et al. (2023) MM’23 53.1 48.2 55.2 62.0 83.8 77.9 82.8 76.7
CCLNet Chen et al. (2023) MM’23 54.0 50.2 56.7 65.1 69.9 65.5 70.2 66.7
PGM Wu and Ye (2023) CVPR’23 57.3 51.8 56.2 62.7 69.5 65.4 69.9 65.2
GUR* Yang et al. (2023c) ICCV’23 61.0 57.0 64.2 69.5 73.9 70.2 75.0 69.9
MMM Shi et al. (2024) arXiv’24 61.6 57.9 64.4 70.4 89.7 80.5 85.8 77.0
PCLMP Ours 64.4 58.7 69.5 74.4 84.3 80.7 82.7 78.4
PCLMP+MMM Enhanced 65.9 61.8 70.3 74.9 89.6 83.7 87.0 80.9
4 Experiment
We conduct extensive experiments to validate the superiority of our proposed method. First, we provide the detailed experiment setting, which contains datasets, evaluation protocols, and implementation details. Then, we compare our method with many state-of-the-art VI-ReID methods and conduct ablation studies. In addition, to better illustrate our method, we also exhibit further analysis. If not specified, we conduct analysis experiments on SYSU-MM01 in the single-shot & all-search mode.
Table 2: Ablation studies on SYSU-MM01 in all search mode and indoor search mode. “Baseline” means the model trained following PGM  Wu and Ye (2023). Rank-R accuracy(%) and mAP(%) are reported.
Method All Search Indoor Search
Index Baseline HPCL DPCL PCL Rank-1 mAP Rank-1 mAP
1 56.3 51.7 60.5 66.2
2 59.1 54.4 63.6 68.8
3 62.1 56.8 65.2 69.8
4 63.7 57.8 67.0 72.6
5 64.4 58.7 69.5 74.4
4.1 Experiment Setting
Dataset. We evaluate our method on two common benchmarks in VI-ReID: SYSU-MM01 Wu et al. (2017) and RegDB Nguyen et al. (2017). SYSU-MM01 is a large-scale public benchmark for the VI-ReID task, which contains 491 identities captured by four RGB cameras and two IR cameras in both outdoor and indoor environments. In this dataset, 22,258 RGB images and 11,909 IR images with 395 identities are collected for training. In the inference stage, the query set consists of 3,803 IR images with 96 identities and the galley set contains 301 randomly selected RGB images. RegDB is collected by an RGB camera and an IR camera, which contains 4,120 RGB images and 4,120 IR images with 412 identities. To be specific, the dataset is randomly divided into two non-overlapping sets: one set is used for training and the other is for testing.
Evaluation Protocols. The experiment follows the standard evaluation settings in VI-ReID, i.e., Cumulative Matching Characteristics (CMC)  Ye et al. (2018) and mean Average Precision (mAP). Specifically, we present the results of all-search and indoor-search modes for SYSU-MM01. The RegDB dataset contains two evaluation settings: Visible2Thermal and Thermal2Visible, i.e., Visible images as query images to search gallery sets consisting of infrared images and vice versa.
Implementation Details. The proposed method is implemented on four NVIDIA 3090 GPUs with PyTorch. We adopt the feature extractor in AGW Ye et al. (2022), which is initialized with ImageNet-pretrained weights to extract 2048-dimensional features. During the training stage, the input images are resized to 288$\times$144. We follow augmentations in  Ye et al. (2021) for data augmentation. In one batch, we randomly sample 16 person IDs, and each person ID samples 16 instances. We set $M$ to be 16 for computational convenience. We adopt Adam as an optimizer to train the model with an initial learning rate of 3.5e-4, decaying to 1/10 of the previous value times every 20 epochs. The number of epochs is 100, in which the first 50 epochs are trained by contrastive loss with the centroid prototype. For the last 50 epochs, we train the model by contrastive loss with both the hard and dynamic prototypes. $E_{CPCL}$ is 50. At the beginning of each epoch, we utilize the DBSCAN Ester et al. (1996) algorithm to generate pseudo labels. During the inference stage, we use the momentum backbone $\phi_{m}$ to extract features and take the features of the global average pooling layer to calculate cosine similarity for retrieval. The momentum value $\alpha$ and $\beta$ is set to 0.1 and 0.999, respectively. The temperature hyper-parameter $\tau$ is set to 0.05 and the weighting hyper-parameter $\lambda$ in Eq.(20) is 0.5.
4.2 Results and Analysis
To comprehensively evaluate our method, we compare our method with 20 supervised VI-ReID methods, 4 semi-supervised VI-ReID methods, and 9 unsupervised VI-ReID methods. The comparison results on the SYSU-MM01 and RegDB are reported in Tab. 1.
Comparison with USVI-ReID Methods. As shown in Tab. 1, our method achieves superior performance compared with state-of-the-art USVI-ReID methods. MMM Shi et al. (2024) is proposed to establish reliable cross-modality correspondences and is also the current best-performing method. Our method with MMM can achieve 65.9% in Rank-1 and 61.8% in mAP, which surpasses that of MMM by a large margin of 4.3% and 3.9%. Notably, our method even without MMM gains the best performance with 64.4% in Rank-1 and 58.7% in mAP. Although existing USVI-ReID methods mentioned in Tab. 1 have made great progress in the USVI-ReID task, the neglect of diverse individual information hinders their further improvement. They overlook diverse individual information, which often constitutes hard samples. Thus, we propose progressive contrastive learning with hard and dynamic prototypes to mine hard samples, which can guide the model to learn more robust and discriminative features.
Comparison with SSVI-ReID Methods. There are four SSVI-ReID methods proposed to alleviate the problem of labeling cost by using a part of data annotations and they have made some progress. However, the application of these methods in real scenarios is still limited due to the fact that data annotations are relatively expensive. Remarkably, our method’s performance without any annotations surpasses that of all methods relying on partial annotation.
Comparison with SVI-ReID Methods. Surprisingly, our method without annotation outperforms several SVI-ReID methods, e.g., DDAG Ye et al. (2020), AGW Ye et al. (2022), NFS Chen et al. (2021), LbA Park et al. (2021). This shows the immense competitiveness of PCLMP compared to SVI-ReID methods that rely on complete data labeling. The superior performance of PCLMP mainly benefits from the hard prototype and dynamic prototype contrastive learning. Additionally, it’s worth acknowledging that there is still a significant disparity between PCLMP and the state-of-the-art fully-supervised results.
Figure 3: (a) The effect of hyper-parameter $\lambda$ with different values. (b) The effect of hyper-parameter $k$ with different values. (c) Comparisons with ARI values of different methods.
Figure 4: The t-SNE visualization of 10 randomly selected identities. Different color indicates different IDs. Circle means visible features and the pentagram means infrared features.
4.3 Ablation Study
In this section, we conduct ablation studies on the SYSU-MM01 dataset in both all-search and indoor-search modes to show the effectiveness of each component in our method. The results are shown in Tab. 2.
Baseline Settings. We use PGM Wu and Ye (2023) as our baseline. Although PGM has achieved a promising performance on the USVI-ReID task, the neglect of hard samples hinders its further improvement.
Effectiveness of HPCL. The HPCL is proposed to mine hard samples at the cluster level. As shown in Tab. 2, When adding the HPCL on Baseline, the performance improves a large margin of 5.8% in Rank-1 and 5.1% in mAP, respectively. It shows that hard samples can be effectively mined using hard prototype contrast learning, facilitating the model to learn more discriminative features.
Effectiveness of DPCL. The DPCL is proposed to mine hard samples at the instance level. The results show that Rank-1 accuracy can be improved by 2.8% in Rank-1 and 2.7% in mAP when adding the DPCL on Baseline, which confirms that contrastive learning with dynamic prototype can further mine hard samples.
Effectiveness of PCL. PCL is introduced to smoothly shift the model’s attention from simple samples to hard ones. The results show that Rank-1 accuracy can be improved by about 1% in Rank-1 and mAP compared to adding simultaneously the HPCL and DPCL on the Baseline. This confirms that progressive contrastive learning plays a valuable role in assisting HPCL and DPCL in mining hard samples.
Surprisingly, contrastive learning with both hard and dynamic prototypes significantly exceeds the baseline by a large margin of 8.1% in Rank-1 and 7.0% in mAP. The HPCL and DPCL can complement each other to mine hard samples, which effectively guides the network to learn more robust and discriminative features.
4.4 Further Analysis
Hyper-parameter analysis. Hyper-parameter $\lambda$ is a weighting parameter to trade-off $L_{HPCL}$ and $L_{DPCL}$. Fig. 3 (a) presents the results under different values of $\lambda$. We can observe that when $\lambda$ is small, i.e., $L_{DPCL}$ contributes more to the model, the performance degrades. However, when $\lambda$ is large, i.e., $L_{HPCL}$ contributes heavily to the model, the model both achieves superior performance. Note that when $\lambda=1$, i.e., the proposed method is trained without DPCL, the performance drops significantly. $\lambda$ is finally set to 0.5 and our method achieves the best performance of 64.4% in Rank-1. Moreover, we also analyze the effect of the number of hard samples at hard prototype. As shown in Fig. 3 (b), we vary the $k$ from 1 to 3 and keep the other hyper-parameters fixed, which shows that PCLMP achieves the best performance when $k=1$.
The ARI metric. Following MMM Shi et al. (2024), we utilize the Adjusted Rand Index (ARI) metric for clustering evaluation. The larger the ARI value, the higher the clustering quality. In Fig. 3 (c), “RGB” and “IR” denote the ARI values of visible and infrared clusterings, which can measure the quality of visible and infrared pseudo-labels. “ALL” means the ARI values of overall clusterings, which can evaluate the reliability of cross-modality correspondences. PCLMP surpasses other methods significantly on all of the mentioned ARI values, which demonstrates our method can effectively mine hard samples to improve clustering quality.
Visualization. As shown in Fig. 4, we visualize the t-SNE map of 10 randomly chosen identities from SYSU-MM01. Compared to the baseline, the distribution of the same identity from the same modality is more compact and the distance of the same identity from different modalities is closer together. Moreover, some hard samples in the baseline are incorrectly clustered, while these hard samples are well clustered in our method, which shows the effectiveness of the proposed PCLMP.
5 Conclusions
In this paper, we propose a novel method for USVI-ReID called Progressive Contrastive Learning with Multi-Prototype (PCLMP), which learns both shared and diverse information. To be specific, we design Hard Prototype Contrastive Learning to mine distinctive yet significant information and Dynamic Prototype Contrastive Learning to preserve intrinsic diversity features. Furthermore, we introduce a progressive learning strategy to incorporate both HPCL and DPCL into the model. Extensive experiments demonstrate that PCLMP outperforms state-of-the-art USVI-ReID methods.